# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ Abstract
We propose a data-driven dynamic factor framework where a response variable $y(t)\in\mathbb{R}^{n}$ depends on a high-dimensional set of covariates $x(t)\in\mathbb{R}^{m}$, without imposing any parametric model on the joint dynamics. Leveraging Anisotropic Diffusion Maps, a nonlinear manifold learning technique introduced by Singer & Coifman, our framework uncovers the joint dynamics of the covariates and responses in a purely data-driven way. We approximate the embedding dynamics using linear diffusions, and exploit Kalman filtering to predict the evolution of the covariates and response variables directly from the diffusion map embedding space. We generalize Singer’s convergence rate analysis of the graph Laplacian from the case of independent uniform samples on a compact manifold to the case of time series arising from Langevin diffusions in Euclidean space. Furthermore, we provide rigorous justification for our procedure by showing the robustness of approximations of the diffusion map coordinates by linear diffusions, and the convergence of ergodic averages under standard spectral assumptions on the underlying dynamics. We apply our method to the stress testing of equity portfolios using a combination of financial and macroeconomic factors from the Federal Reserve’s supervisory scenarios. We demonstrate that our data-driven stress testing method outperforms standard scenario analysis and Principal Component Analysis benchmarks through historical backtests spanning three major financial crises, achieving reductions in mean absolute error of up to 55% and 39% for scenario-based portfolio return prediction, respectively.
@@ 1 Introduction
The emerging field of data-driven dynamical systems aims to learn the potentially high-dimensional and nonlinear dynamical description of a data set without parametric assumptions. Some of the primary goals in this area include finding stochastic differential equation (SDE) descriptions of discretely sampled time series data, extracting interpretable stochastic models from experimental data, and finding simple linear descriptions for high-dimensional and nonlinear dynamical systems. Herein, we propose a data-driven approach to the problem of dynamic factor modeling in the supervised learning setting where a response variable $y(t)\in\mathbb{R}^{n}$ depends on a high-dimensional set of covariates $x(t)\in\mathbb{R}^{m}$. Dynamic factor models typically find a state-space representation of a set of observed measurements, and operate under the premise that a few latent dynamic factors drive the co-movements of a multivariate time series. These latent factors themselves follow a time series process, commonly a vector auto-regressive (VAR) process, and restrictive assumptions are often imposed. (For an in-depth survey of classical dynamic factor models, we recommend [StockWatson2016].) In this work, we propose a data-driven dynamic factor framework which accounts for nonlinear joint covariate and response dynamics.
Standard Dynamic Factor Models usually uncover the latent factors from a single set of measurements $x(t)$. However, if one wants to use these latent factors as covariates under a supervised learning framework where $y(t)=f(x(t))+\varepsilon_{t}$, these models do not jointly consider the variation of $y(t)$ when uncovering the dynamic latent factors. As a consequence, the estimated sparse set of latent factors may not contain explanatory power for the responses. Rather than assuming a parametric model for the covariate dynamics, we leverage the geometry of the time series data and exploit Anisotropic Diffusion Maps (ADM), a nonlinear dimensionality reduction technique introduced in [singer2008nonlinear]. This allows us to learn the intrinsic dynamics of the high-dimensional covariate vector via lower-dimensional diffusion map embeddings, while jointly preserving much of the explanatory power of the time series responses. We define a state space completely in terms of the diffusion map embeddings that allows us to approximate the dynamics of the system using Kalman filtering, and perform response prediction and covariate sampling directly from the embedding space.
@@
We treat the observation vector ${(x(t),y(t))\in\mathbb{R}^{m+n}}$ as the state of a dynamical system at time $t$. We assume the covariates and response variables are nonlinear functions of a true, possibly lower-dimensional, underlying state of the system $\theta_{t}\in\mathbb{R}^{d}$, which is assumed to satisfy a Langevin diffusion. These nonlinear functions restrict the trajectories of the observations to lie on some (unknown) manifold $\mathcal{M}$. Figure 1 displays a broad overview of our setting. Our framework relies on three key steps: (1) uncover latent dynamic factors, (2) estimate underlying dynamics from these factors, and (3) predict responses from learned functional relationships.
Figure 1: A diagram of the dynamical system setting.
Our framework relies on constructing dynamic diffusion coordinate embeddings of the observations that serve as proxies for the dynamics of $\theta$, combining state space modeling with diffusion maps. Commonly, dimensionality reduction methods, when applied to time series data, do not explicitly consider the temporal component in the data. This can be limiting in many applications, for instance, when dealing with high-dimensional financial market data where a portfolio manager (PM) might want to study how shocks to a given market factor propagate through time. By integrating Kalman filtering techniques with diffusion maps and extending the framework first introduced in [shnitzer2020diffusion], we construct dynamic diffusion coordinate embeddings which explicitly factor in the temporal component of the data. We construct our state space equations by leveraging a linear lifting operator that takes diffusion map embeddings back into the original measurement space.
@@
The central mathematical contributions of our work are as follows: Applying concentration inequalities for Markov processes, in Theorem 4.3 we generalize a result on the convergence of graph Laplacians from [singer_graph_2006] to the time series case, and show that the eigenfunctions of the Fokker–Planck operator associated with $\theta$ can be learned. We provide rigorous justification for reconstructing points in the original data space via the lifting operator by establishing its convergence to an $\mathbb{L}^{2}$-projection with respect to the invariant measure of $\theta$ (see Proposition 4.2). The rationale for utilizing diffusion maps and defining a state space framework completely in terms of these embeddings is the observation, first mentioned in [coifman2008diffusion], that the dynamics of the diffusion coordinate embeddings are approximately linear. Under standard spectral assumptions on the generator of $\theta$, we show in Subsection 4.3 that the approximation error using linear dynamics remains small when averaging over the data.
Since we approximate the intrinsic dynamics with a linear SDE in terms of the diffusion coordinate embeddings, our framework does not require us to discover the explicit joint dynamics of the observations $(x(t),y(t))$. Leveraging linear state-space equations, we devise a procedure to sample dynamic embeddings while conditioning on an information set defined in the original measurement space. It is well known that dimensionality reduction techniques give rise to the challenge of interpretability of the embeddings. In a framework such as ours that takes place completely in the embedding space, it is hence desirable to define fully interpretable conditioning events in the original measurement space. To define such a sampling procedure, we leverage the power of the linear lifting operator in a three step process: (1) lift statistics from the embedding space to the original measurement space, (2) condition on the desired information set in the covariate space and obtain the conditional distribution in terms of $(x(t),y(t))$, (3) restrict back to the embedding space with a pseudoinverse of the lifting operator. Our sampling procedure simultaneously learns the intrinsic geometry and dynamics from the data, and then generates sample paths for the diffusion coordinate embeddings and by extension of $(x(t),y(t))$. Of course, it is known that dimensionality reduction techniques do not necessarily preserve statistics in the data, hence approximating statistics in the measurement space with their lifted counterparts incurs some error. Our robustness results presented in Section 4 provide theoretical support for the applicability of our method.
We demonstrate the power of our framework for applications, focusing on problems from mathematical finance, where dynamical systems are ubiquitous. Continuous time processes, typically represented by SDEs, include interest rate models such as the Vasicek process, credit risk models such as Cox–Ingersoll–Ross, and derivatives pricing models such as Black–Scholes–Merton. Discrete time models are also widely present in financial applications; these include models for stochastic volatility such as ARCH and GARCH, and models for returns such as the standard autoregressive process. Strong parametric and modeling assumptions are often applied when fitting real-world data with stochastic processes.
@@
Under comparably weak assumptions, our approach is applicable to a wide variety of problems. These include, for instance, asset pricing, where often one works with a high-dimensional “zoo” of candidate factors to explain the cross-section of stock returns. It is often desirable to construct a sparse set of factors, often latent, from the factor zoo in order to construct a model for the cross-section of returns. Our approach is well-suited for such a task as it produces lower-dimensional dynamic embeddings which propagate linearly, and also possess explanatory power comparable to the original high-dimensional set of factors. Another application is nowcasting (see [giannone_nowcasting_2008]), where one makes use of higher frequency/quality data to improve predictions on lower frequency/quality data. A third key application, with which we validate our framework via an empirical investigation, is financial stress testing and scenario analysis, which deals with predicting an asset or portfolio’s evolution conditional on a set of shocks to the risk factors that drive their value (see, for instance, [haugh2020scenario]). Within the context of stress testing equity portfolios, we demonstrate the superiority of our framework over several comparison benchmarks including Principal Component Analysis (PCA) via a series of historical rolling backtests. Our historical backtests span three major financial crisis periods and we demonstrate reductions in mean absolute error (MAE) of up to 55% and superior accuracy up to 77.78% of the time when utilizing our framework over the benchmarks.
@@
The paper is organized as follows. Section 2 relates our work to the literature. Section 3 introduces our dynamic factor modeling framework and conditional sampling procedure. Section 4 presents key theoretical results, including convergence analysis of the graph Laplacian and the robustness of approximating the diffusion coordinates by linear diffusions. Section 5 describes the application of our method to financial stress testing: in 5.1 we describe the comparison benchmarks and 5.2 defines the historical backtesting procedure. Section 6 describes the data and the empirical results. We conclude in Section 7. Appendices A, B, C, D and E provide further details on the data, empirical results, implementation of benchmarks, and a lengthy calculation used for the analysis of the graph Laplacian.

### v2

@@ Abstract
We introduce a data-driven dynamic factor framework for modeling the joint evolution of high-dimensional covariates and responses without parametric assumptions. Standard factor models applied to covariates alone often lose explanatory power for responses. Our approach uses anisotropic diffusion maps, a manifold learning technique, to learn low-dimensional embeddings that preserve both the intrinsic geometry of the covariates and the predictive relationship with responses. For time series arising from Langevin diffusions in Euclidean space, we show that the associated graph Laplacian converges to the generator of the underlying diffusion. We further establish a bound on the approximation error between the diffusion map coordinates and linear diffusion processes, and we show that ergodic averages in the embedding space converge under standard spectral assumptions. These results justify using Kalman filtering in diffusion-map coordinates for predicting joint covariate-response evolution. We apply this methodology to equity-portfolio stress testing using macroeconomic and financial variables from Federal Reserve supervisory scenarios, achieving mean absolute error improvements of up to 55% over classical scenario analysis and 39% over principal component analysis benchmarks.
@@ 1 Introduction
The emerging field of data-driven dynamical systems aims to learn the potentially high-dimensional and nonlinear dynamical description of a data set without parametric assumptions. Some of the primary goals in this area include finding stochastic differential equation (SDE) descriptions of discretely sampled time series data, extracting interpretable stochastic models from experimental data, and finding simple linear descriptions for high-dimensional and nonlinear dynamical systems. Herein, we propose a data-driven approach to the problem of dynamic factor modeling in the supervised learning setting where a response variable $y(t)\in\mathbb{R}^{n}$ depends on a high-dimensional set of covariates $x(t)\in\mathbb{R}^{m}$. A key challenge in this setting is that dimensionality reduction applied to covariates alone, such as principal component analysis, may discard information critical for predicting responses, even when the intrinsic geometry of $x(t)$ is preserved. Our framework addresses this by learning embeddings jointly from $(x(t),y(t))$ that capture both the dynamics of the covariates and their relationship to the responses.
@@
We treat the observation vector $z(t)={(x(t),y(t))\in\mathbb{R}^{m+n}}$ as the state of a dynamical system at time $t$. We assume the covariates and response variables are nonlinear functions of a true, possibly lower-dimensional, underlying state of the system $\theta_{t}\in\mathbb{R}^{d}$, which is assumed to satisfy a Langevin diffusion. These nonlinear functions restrict the trajectories of the observations to lie on some (unknown) manifold $\mathcal{M}$.
To fix the mathematical setting, consider an $\mathbb{R}^{d}$-valued stochastic process $\theta=(\theta_{t})_{t\geq 0}$ and an $\mathbb{R}^{m+n}$-valued process $Z=(Z_{t})_{t\geq 0}$ satisfying the equations
$\mathrm{d}\theta_{t}$ $=-\nabla U(\theta_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t},$ (1.1)
$Z_{t}$ $=H(\theta_{t}),$ (1.2)
where $W$ is a $d$-dimensional standard Brownian motion. The potential $U:\mathbb{R}^{d}\to\mathbb{R}$ is assumed to grow fast enough so that $\theta$ admits the invariant measure $\mu(\mathrm{d}w)\propto e^{-U(w)}\mathrm{d}w$, and the infinitesimal generator of $\theta$, given by $\mathscr{L}:=\Delta-\nabla U\cdot\nabla$, possesses a discrete spectrum. (We restate our assumptions formally in Section 3.1 below.) The function $H:\mathbb{R}^{d}\to\mathbb{R}^{m+n}$} is assumed to be $C^{2}$-smooth and $\mathbb{L}^{2}(\mu)$-integrable. We view these equations as a degenerate hidden Markov model (HMM): $Z$ is an observed process which depends (deterministically) on the evolution of the latent state $\theta$, which in turn evolves stochastically according to a Langevin diffusion. Given $N+1$ discrete observations of a single $Z$-trajectory, $\{Z(t_{i})\}_{i=0}^{N}$, and without any knowledge of $H,U,$ or even the latent dimension $d$, we wish to learn the dynamics of $Z$. That is, we wish to build an accurate data-driven approximation to carry out tasks such as prediction, conditional sampling, interpolation, and generation of new sample trajectories.
In broad terms, we adopt a spectral approach, whereby $H$ is approximated using $\ell\geq 1$ principal eigenfunctions $(\varphi_{1},\dots,\varphi_{\ell})$ of the infinitesimal generator $\mathscr{L}$. As $H\in\mathbb{L}^{2}(\mu)$, we have that
$Z_{t}=H(\theta_{t})\approx\mathbf{H}(\varphi_{1}(\theta_{t}),\dots,\varphi_{\ell}(\theta_{t}))^{\top},$ (1.3)
for some matrix $\mathbf{H}\in\mathbb{R}^{(m+n)\times\ell}$. If $\varphi_{i}$ is associated with eigenvalue $\lambda_{i}<0$, then Itô’s formula yields
$\mathrm{d}\varphi_{i}(\theta_{t})=-\lambda_{i}\varphi_{i}(\theta_{t})\mathrm{d}t+\sqrt{2}\nabla\varphi_{i}(\theta_{t})\cdot\mathrm{d}W_{t}.$ (1.4)
Approximating $(\varphi_{1}(\theta_{t}),\dots,\varphi_{\ell}(\theta_{t}))^{\top}$ by a linear diffusion $\xi=(\xi_{t})_{t\geq 0}$ in $\mathbb{R}^{\ell}$, we arrive at a linear HMM:
$\mathrm{d}\xi_{t}$ $=-\Lambda\xi_{t}\mathrm{d}t+\Gamma\mathrm{d}B_{t},$ (1.5)
$Z_{t}$ $=\mathbf{H}\xi_{t}+\eta_{t},$ (1.6)
where $\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{\ell})$, $\Gamma\in\mathbb{R}^{\ell\times\ell}$, and $B$ is a standard $\ell$-dimensional Brownian motion. Here, $\eta=(\eta_{t})_{t\geq 0}$ is a noise term incorporating the approximation error between $\xi$ and the eigenfunctions, as well as the truncation error from the expansion of $H$.
The learning problem in this setting amounts to estimating the eigenvalues which comprise $\Lambda$ and the projection coefficients $\mathbf{H}$. Assuming a Gaussian model for $(\eta_{t})_{t\geq 0}$, we arrive at a linear-Gaussian state-space model for $Z$. Using standard Kalman filtering techniques, one may fit $\Gamma$ and carry out prediction, conditional sampling, and so on. We view this non-parametric approach, introduced as the diffusion Kalman filter (DKF) and applied to object tracking in [36], as an alternative to the extended Kalman filter (see, for instance, [25]). In Subsection 5, we generalize this methodology to incorporate both the covariates $x(t)$ and response variables $y(t)$, yielding the joint diffusion Kalman filter (JDKF).
As a special case, assuming a linear relationship between $x(t)$ and $y(t)$ yields a data-driven approach to linear multi-factor modeling, which can be compared and interpreted alongside standard linear models such as principal component analysis. By contrast, standard dynamic factor models typically uncover latent factors from a single set of measurements $x(t)$. In supervised settings where $y(t)$ is a noisy function of $x(t)$, factors learned from $x(t)$ alone may fail to retain explanatory power for $y(t)$. Motivated by this limitation, our approach learns diffusion-coordinate embeddings jointly from $(x(t),y(t))$, and models their evolution through a linear state-space approximation. This ensures that the learned factors retain explanatory power for $y(t)$ while still capturing the intrinsic dynamics of $x(t)$. The key insight is to treat both covariates and responses as noisy observations of a common latent state $\theta$ evolving according to unknown dynamics, and to learn this latent evolution in a data-driven way via diffusion maps. This yields a supervised dynamic factor framework that supports prediction and conditional sampling, and is formalized in Section 5.
@@
Our aims in this work are threefold: (1) we show that $\Lambda$ and $\mathbf{H}$ can be learned using anisotropic diffusion maps and provide a quantitative error bound, (2) we justify the reduction to a linear HMM by bounding the approximation error between the eigenfunctions and linear diffusions, and (3) we demonstrate the effectiveness of our approach for dynamic factor modeling with an application to conditional sampling in the context of the stress testing of equity portfolios. The mathematical contributions of our work are as follows: applying concentration inequalities for Markov processes, in Theorem 4.3 we generalize a result on the convergence of graph Laplacians from [38] to the time series case, showing that the eigenfunctions of $\mathscr{L}$ can be learned. With Proposition 4.2, we provide rigorous justification for reconstructing points in the original data space via the lifting operator by establishing its convergence to an $\mathbb{L}^{2}$-projection with respect to the invariant measure of $\theta$. The rationale for utilizing diffusion maps and defining a state space framework completely in terms of these embeddings is the observation, first mentioned in [13], that the dynamics of the diffusion coordinate embeddings are approximately linear. Under standard spectral assumptions on the generator of $\theta$, we show in Subsection 4.3 that the approximation error using linear dynamics remains small when averaging over the data.
1.1 Manifold Learning with Diffusion Maps
We carry out the learning of $\Lambda$ and $\mathbf{H}$ using the diffusion maps algorithm, which we now outline following [3, 41]. Consider first the case where $H=I$ so that $Z=\theta$, and we wish to determine the eigenvalues of the infinitesimal generator $\mathscr{L}$ from samples $\{\theta(t_{i})\}_{i=0}^{N}$. For $\varepsilon>0$, let $P_{\varepsilon}=\exp(\varepsilon\mathscr{L})$ denote the Markov transition operator associated with $\theta$, and let $p_{\varepsilon}$ denote the associated transition kernel with respect to the invariant measure $\mu$. As noticed in [3], $P_{\varepsilon}$ and $\mathscr{L}$ share eigenfunctions, and the eigenvalues of $P_{\varepsilon}$ can be mapped to those of -$\mathscr{L}$ by $\lambda\mapsto-\frac{1}{\varepsilon}\log(\lambda)$. One may approximate $P_{\varepsilon}$ by the random operator $P_{\varepsilon}^{N}$ which acts on continuous bounded functions $f:\mathbb{R}^{d}\to\mathbb{R}$ by
$P_{\varepsilon}^{N}f(w):=\frac{1}{N+1}\sum_{i=0}^{N}p_{\varepsilon}(w,\theta(t_{i}))f(\theta(t_{i})).$ (1.7)
This can be rewritten as
$P_{\varepsilon}^{N}f(w)=\int p_{\varepsilon}(w,y)f(y)\mu_{N}(\mathrm{d}y)$ (1.8)
where $\mu_{N}=\frac{1}{N+1}\sum_{i=0}^{N}\delta_{\theta(t_{i})}$ denotes the empirical measure defined by the data. As $N\to\infty$, one hopes that $\mu_{N}$ approximates the invariant measure $\mu$, that $P_{\varepsilon}^{N}\to P_{\varepsilon}$, and the eigenvalues/functions/spaces of $P_{\varepsilon}^{N}$ converge to those of $P_{\varepsilon}$.
The construction of $P_{\varepsilon}^{N}$ requires a priori knowledge of the transition kernel $p_{\varepsilon}$. To overcome this obstruction, we approximate $p_{\varepsilon}$ using a Gaussian kernel
$k_{\varepsilon}(w,y)=\exp(-\|w-y\|^{2}/4\varepsilon).$ (1.9)
This is justified since the short-time asymptotic for the heat kernel of a Langevin diffusion satisfies
$\lim_{t\downarrow 0}4\varepsilon\log(p_{\varepsilon}(w,y))=-\|w-y\|^{2}.$ (1.10)
(this follows from applying a Girsanov transformation; and see also [23], for background). Consider the operator $T_{\varepsilon}^{N}$ which acts on bounded continuous $f$ by
$T_{\varepsilon}^{N}f(w):=\frac{\sum_{i=0}^{N}k_{\varepsilon}(w,\theta(t_{i}))f(\theta(t_{i}))}{\sum_{i=0}^{N}k_{\varepsilon}(w,\theta(t_{i}))}.$ (1.11)
Setting
$d_{\varepsilon}^{N}(w):=\sum_{i=0}^{N}k_{\varepsilon}(w,\theta(t_{i}))=\int k_{\varepsilon}(w,y)\mu_{N}(\mathrm{d}y),$ (1.12)
we may rewrite $T_{\varepsilon}^{N}f(w)$ as
$T_{\varepsilon}^{N}f(w)=\frac{1}{d_{\varepsilon}^{N}(w)}\int k_{\varepsilon}(w,y)f(y)\mu_{N}(\mathrm{d}y).$ (1.13)
Supposing that $T_{\varepsilon}^{N}$ converges to some limiting $T_{\varepsilon}$, and $T_{\varepsilon}$ is close to $P_{\varepsilon}$ for small $\varepsilon$, then the eigenvalues/functions/spaces of $P_{\varepsilon}$ can be approximated by those of $T_{\varepsilon}^{N}$ for large $N$. Furthermore, the spectrum and eigenfunctions of $T_{\varepsilon}^{N}$ can be computed explicitly as they are in a one-to-one correspondence with the matrix $\mathbf{P}$ which has entries
$\mathbf{P}_{ij}=\frac{k_{\varepsilon}(\theta(t_{i}),\theta(t_{j}))}{d_{\varepsilon}^{N}(\theta(t_{i}))},\quad i,j=1,\dots,N,$
see [41, Proposition 7 ].
[…]
$\|w-\widetilde{w}\|^{2}$ $=\frac{1}{2}(z-\widetilde{z})^{T}\left((J_{H}J_{H}^{T})^{-1}(z)+(J_{H}J_{H}^{T})^{-1}(\widetilde{z})\right)(z-\widetilde{z})+O(\|z-\widetilde{z}\|^{4})$ (1.14)
Although $H$ is not observed, the Jacobian $J_{H}$ can be deduced from the quadratic covariation process of $Z$. To wit, applying Itô’s formula to a component $Z^{i}=H^{i}(\theta)$ gives
$\mathrm{d}Z_{t}^{i}=\mathscr{L}H_{i}(\theta_{t})\mathrm{d}t+\sqrt{2}\nabla H_{i}(\theta_{t})\cdot\mathrm{d}W_{t}$
and hence
$\mathrm{d}\langle Z^{i},Z^{j}\rangle_{t}=2\nabla H_{i}(\theta_{t})\cdot\nabla H_{j}(\theta_{t})\mathrm{d}t,\quad i,j=1,\dots,m+n.$
Therefore, $\|\theta(t_{i})-\theta(t_{j})\|^{2}$ can be approximated by the squared modified Mahalanobis distance between samples of $Z$:
$d(Z(t_{i}),Z(t_{j}))^{2}:=\frac{1}{2}(Z(t_{i})-Z(t_{i}))^{T}(C^{-1}(t_{i})+C^{-1}(t_{j}))(Z(t_{i})-Z(t_{j})),$
where $C$ is an estimator for the instantaneous quadratic covariation process of $Z$.
1.2 Applications
We demonstrate the applicability of our framework in problems from mathematical finance, where stochastic dynamical systems play a central role. Classical continuous and discrete-time models, including interest rate, credit risk, volatility, and return models, are typically specified parametrically and tailored to low-dimensional settings. Our data-driven approach is designed to operate under substantially weaker modeling assumptions and to scale naturally to high-dimensional environments.
@@ 1.2 Applications
We validate our framework in the context of financial stress testing and scenario analysis through a detailed empirical investigation. These settings involve predicting the evolution of an asset or portfolio conditional on shocks to the risk factors that drive its value (see, for instance, [22]). Within the context of stress testing of equity portfolios, we demonstrate the superiority of our framework over several comparison benchmarks including Principal Component Analysis (PCA) via a series of historical rolling backtests. Our historical backtests span three major financial crisis periods and we demonstrate reductions in mean absolute error (MAE) of up to 55% and superior accuracy up to 77.78% of the time when utilizing our framework over the benchmarks.
@@
The rest of the paper is organized as follows. Section 2 relates our work to the literature. Section 3 introduces some preliminary concepts and tools necessary for our framework, including the dynamical system setting and the diffusion map construction method. Section 4 presents key theoretical results, including convergence analysis of the graph Laplacian and the robustness of approximating the diffusion coordinates by linear diffusions. Section 5 introduces our dynamic factor modeling framework and conditional sampling procedure. Section 6 describes the application of our method to financial stress testing: in 6.1 we describe the comparison benchmarks and 6.2 defines the historical backtesting procedure. Section 7 describes the data and the empirical results. We conclude in Section 8. The appendices provide further details on the data, empirical results, implementation of benchmarks, and a lengthy calculation used for the analysis of the graph Laplacian.